Four oranges and five apples cost $2.00. Three oranges and four apples cost $1.56. Find the cost of an orange and the cost of an apple.
Answer by stanbon(75887)

Four oranges and five apples cost $2.00. Three oranges and four apples cost $1.56. Find the cost of an orange and the cost of an apple.
--------------------
4r + 5a = 200 cents
3r + 4a = 156 cents
-------------------------
Modify to prepare for elimination:
12r + 15a = 600
12r + 16a = 624
------------------------
Subtract 1st from 2nd to solve for "a":
a = 24 cents (price of an apple)
-----
Substitute to solve for "r":
4r + 5*24 = 200
4r + 120 = 200
4r = 80
r = 20 cents (price of an orange)
======================================
Cheers,
Stan H.